OpenStudy (anonymous):

Sum of all numbers from 100 to 999, divisible by 11

OpenStudy (zarkon):

@dpasingh 880/11=80

OpenStudy (anonymous):

Oh yes sorry

OpenStudy (zarkon):

\[81\times\frac{110+990}{2}=44550\]

OpenStudy (anonymous):

The numbers from 100 to 999 divisible by 11 are as given below.. 110, 121, 132,.......990 Therefore here first term a= 110, commomn difference d=11 and tn =990 Now we shall find the number of terms n since tn = a+(n-1)d i.e. 990= 110+(n-1)11 i.e. 990- 110=(n-1)11 i.e. 880=(n-1)11 i.e. 80=n-1 n= 81 Now sum of given numbers is given by Sn = n/2[a+tn] i.e. S81= 81/2[110+990]= 81/2[1100]=81*550=44550 i.e. S81=44550

OpenStudy (anonymous):

Tyvm :)